# Chiral order emergence driven by quenched disorder

### v2

@@ Abstract
For $0.05\leq p\leq 0.1$, $\Sigma$ collapses to zero at all temperatures as $L$ increases. This behavior becomes more pronounced as $p$ increases[40]. In the thermodynamic limit, we expect the $K_{4}$ order to vanish, which may indicate that site dilution finally acts as a random field and kills the transition for $p>0.05$. At low temperatures, the chiral order is present below a critical temperature that increases with the dilution, but saturates at $p\simeq 0.1$ (see Fig. 1). For $0.1\leq p\leq 0.2$, the $K_{4}$ order is clearly absent, and the chiral order weakens as $p$ increases with a progressive decrease of the critical temperature. Nevertheless, one observes a significant versatility of behaviors: for $p=[0.14,0.16]$, $C_{V}$ displays a small but noticeable divergence with the system size[40]. Moreover, the maximum of $S_{\chi}$ still diverges with $L$, even though its value decreases due to the site dilution, which isolates some of the complete triangles where $\chi$ is defined. FSSA unveils that this maximum scales as $L^{1.7}$ which is compatible with the Ising universality class.
@@
In summary, we have shown on the example of the $J_{1}J_{3}$ model with magnetic site vacancies that quenched disorder can lead to the emergence of a specific order (here chiral), that was forbidden in the clean phase (with collinear $K_{4}$ order). This phenomenon can be designated as the order by quench disorder mechanism that differs from the usual thermal or quantum order by disorder. Although this expression has been previously employed[6], our proposal is distinct in that it involves the breaking of time-reversal symmetry in the absence of a magnetic field and in contrast to the pure model. A similar birth of chirality is expected for other models such as the triangular or honeycomb lattices with stripe orders[44, 34], due to the lattice decoupling in more than two sublattices (similar to what occurs in the $J_{1}J_{3}$ model). The exploration of classical spin liquids[7] represents another direction for further research on order from quenched disorder.

### v3

@@ Abstract
For $0.05\leq p\leq 0.1$, $\Sigma$ collapses to zero at all temperatures as $L$ increases[40]. This behavior becomes more pronounced as $p$ increases, illustrating that site dilution effectively acts as a random field for the $K_{4}$ order and kills the transition. At low temperatures, the chiral order is present below a critical temperature that increases with the dilution, but saturates at $p\simeq 0.1$ (see Fig. 1). For $0.1\leq p\leq 0.2$, the $K_{4}$ order is clearly absent, and the chiral order weakens as $p$ increases with a progressive decrease of the critical temperature. $C_{V}$ displays a small but noticeable increase with the system size[40]. Moreover, the maximum of $\mathcal{S}_{\chi}$ still diverges with $L$, even though its value decreases due to the site dilution, which isolates some of the complete triangles where $\chi$ is defined. FSSA unveils that this maximum scales as $L^{1.75}$ which is compatible with the Ising universality class (see Fig. 6 right).
@@
In summary, we have shown on the example of the $J_{1}J_{3}$HK model with magnetic site vacancies that quenched disorder can lead to the emergence of a specific order (here chiral) in $O(3)$-symmetric model, that was forbidden in the clean phase (with collinear $K_{4}$ order) and previously only observed in $XY$ models[6, 25]. This phenomenon can be designated as the order by quench disorder mechanism that differs from the usual thermal or quantum order-by-disorder. Although this expression has been previously employed[24], our proposal is distinct in that it involves the breaking of time-reversal symmetry in the absence of a magnetic field and in contrast to the pure model. A similar birth of chirality is expected for other models such as the triangular or honeycomb lattices with stripe orders[44, 34], due to the lattice decoupling in more than two sublattices (similar to what occurs in the $J_{1}J_{3}$HK model). The exploration of the effect of quenched disorder on classical spin liquids[3, 45, 21, 46, 47, 48, 25, 36] represents another direction for further research.